OpenStudy (funnyguy2):

3 - x/9 = 1/4 they are fractions

OpenStudy (funnyguy2):

@TheSmartOne

OpenStudy (funnyguy2):

@King.Void.

OpenStudy (funnyguy2):

@TheSmartOne

OpenStudy (misty1212):

wrong section

OpenStudy (funnyguy2):

this is math

OpenStudy (funnyguy2):

@misty1212

OpenStudy (misty1212):

\[\frac{3-x}{9}=\frac{1}{4}\]? or \[3-\frac{x}{9}=\frac{1}{4}\]?

OpenStudy (funnyguy2):

the first one

OpenStudy (funnyguy2):

@campbell_st

OpenStudy (campbell_st):

cross multiply \[4( 3 - x) = 9 \times 1\] which becomes 12 - 4x = 9 now solve for x

OpenStudy (funnyguy2):

3/4

OpenStudy (campbell_st):

great.